# Thermal Imaging and Radar for Remote Sleep Monitoring of Breathing and Apnea

### v1

@@ II-C Automatic Apnea Detection
While airflow and respiratory effort methods can provide good accuracy in remote apnea detection, an important vital sign that is missed by these methods is SpO2. For non-contact methods, the most common method to achieve remote SpO2 extraction is through two NIR cameras with different wavelength sensitivities. Typically, these methods first detect a remote-photoplethysmography signal [79, 82, 15, 61, 47, 17, 14], followed by application of the ratio-of-ratios method [12, 58]. This general framework has been applied in several instances [50, 80, 75] with various algorithmic innovations to extract a better remote photoplethysmography signal. Other work also detects SpO2 using spectroscopic methods with a multi-aperture camera [69, 70, 39].
@@ Acknowledgment
The authors would like to thank Dr. Jean-Paul Chretien and Dr. Rebecca McFarland for their valuable feedback and comments throughout the course of the project. The authors would also like to thank UCLA Sleep Disorders Center and its staff, especially Julie Toomey and Weiguang Zhong; undergraduate students Rishabh Sharma, Rui Ma, Jianchong Ma, and Julia Craciun for their assistance in developing parts of the preprocessing code; Junaid Ahmad for his work designing protective circuits for data collection; medical student Clare Moffatt for help with data collection; and Dr. Laleh Jalilian for valuable advice in collecting clinical data. Authors on this research were supported by a DARPA Young Faculty Award; Achuta Kadambi was supported by an National Science Foundation (NSF) CAREER award IIS-2046737, Army Young Investigator Program Award, and Defense Advanced Research Projects Agency (DARPA) Young Faculty Award.

### v2

@@ II-C Automatic Apnea Detection
Non-contact methods have been developed to remotely measure SpO2. The most common method to achieve remote SpO2 extraction is through two NIR cameras with different wavelength sensitivities. Typically, these methods first detect a remote-photoplethysmography signal [79, 82, 15, 61, 47, 17, 14], followed by application of the ratio-of-ratios method [12, 58]. This general framework has been applied in several instances [50, 80, 75] with various algorithmic innovations to extract a better remote photoplethysmography signal. Other work also detects SpO2 using spectroscopic methods with a multi-aperture camera [69, 70, 39]. While SpO2 is an important vital sign for sleep monitoring, we determined that we would require custom ground truth SpO2 measurement hardware to implement a noncontact method. This would have been incompatible with our data collection procedures as it would require our hardware to contact the patients.
@@ Acknowledgment
The authors would like to thank UCLA Sleep Disorders Center and its staff, especially Julie Toomey and Weiguang Zhong; undergraduate students Rishabh Sharma, Rui Ma, Jianchong Ma, and Julia Craciun for their assistance in developing parts of the preprocessing code; Junaid Ahmad for his work designing protective circuits for data collection; medical student Clare Moffatt for help with data collection; and Dr. Laleh Jalilian for valuable advice in collecting clinical data. Authors on this research were supported by a DARPA Young Faculty Award; Achuta Kadambi was supported by a National Science Foundation (NSF) CAREER award IIS-2046737, an Army Young Investigator Program Award, and a Defense Advanced Research Projects Agency (DARPA) Young Faculty Award.